# Cyber security of OT networks: A tutorial and overview

### v1

@@ Abstract
Cyber security, Operational Technology
@@ 1 Introduction
Operational Technology (OT) networks are critical infrastructure systems responsible for controlling physical processes in industries such as manufacturing, energy, transportation, and utilities. These networks integrate hardware and software systems, including SCADA (Supervisory Control and Data Acquisition), PLCs (Programmable Logic Controllers), and DCS (Distributed Control Systems), to ensure the smooth operation of industrial machinery and processes. However, with the convergence of OT and IT systems due to increased digitalization and Industry 4.0 initiatives, OT networks have become a prime target for cyber threats in recent time [1] [2] [3]. Unlike traditional IT systems, OT networks face unique cyber-security challenges due to their legacy systems, real-time operational requirements, and the high cost of downtime or failure. Cyber security in OT networks, therefore, focuses on safeguarding critical infrastructure from threats such as malware, ransomware, and state-sponsored attacks, ensuring the continuity, safety, and reliability of essential operational services [4] [5]. In this manuscript, we focus on understanding the key components in OT networks, how OT and IT overlap creates the security considerations for IT-OT Systems where some security concerns originate in IT networks and lateral move to OT networks and other security concerns in OT networks, the attack vectors, some recent OT attacks and its mode of operation and tools is used in modern time in various aspects of cybersecurity. We will also discuss the regulatory framework and standards in OT specific cyber security, emerging trends by adapting new technology in OT secuity space, and finally evaluates its quantitative impact.
@@ 2 OT and its components
Operation Technology (OT) refers to hardware and software systems used for actuation, sensing and monitoring, computing and control of physical devices, processes, and industrial operations. In this manuscript, the context of OT is in critical sectors such as manufacturing, energy, utilities, transportation, and healthcare etc. The key components and systems in modern OT systems (also called Industrial Control System ICS) are as below:
@@ 2.1 Supervisory Control and Data Acquisition (SCADA)
SCADA system provides centralized monitoring and control of complex processes spread across large areas. SCADA systems gather real-time data from remote locations from sensors and instruments located at remote sites [6] and transmit it to central computers and provides technology to control, and visualization as well. It is widely used in industries like power generation, oil exploration, water management, manufacturing etc. The key components of SCADA systems are field instruments and sensors, Remote Terminal Units (RTUs), Programmable Logic Controllers (PLCs), communication infrastructure, SCADA master station, and data historian. Field instruments and sensors are devices that measure various parameters like temperature, pressure, flow, rpm etc in the industrial process. Remote Terminal Units (RTUs) are microprocessor controlled devices that interface with sensors and convert their signals to digital data [7], which is then transmitted to the central system. Programmable Logic Controllers (PLCs) are industrial digital computers that control the process. communication infrastructure consist of the physical network and protocols that transmits data between the field devices and the control systems. It can use various communication methods and protocols, including wired (Ethernet) and wireless (radio, cellular) networks. SCADA master station is the central control system that collects data from RTUs and PLCs, processes it, and presents it to human operators. It also includes Human-Machine Interface (HMI) software for visualization and high level supervisory control. Data Historian is a system for collecting and storing historical data from the SCADA system. This data is used for trend analysis, reporting, and auditing. Details of the components are provided below:
  1. 1.
    Sensors: Sensors gather real-time data on various process parameters such as temperature, pressure, flow, level, and chemical composition etc. It also provide continuous monitoring of the industrial processes to ensure they are operating within specified parameters. By sensing, it enable to introduce automated control by providing feedback to the SCADA system, which can then make adjustments to maintain optimal operation. It also used to detect abnormal conditions and trigger alarms or safety shutdowns to prevent accidents or equipment damage.
  2. 2.
    RTU: A Remote Terminal Unit (RTU) is an electronic device used in industrial and remote monitoring applications to collect data from sensors, process it, and transmit it to a central control system, such as a SCADA (Supervisory Control and Data Acquisition) system. RTUs are essential for monitoring and controlling equipment and processes in geographically dispersed locations. RTU is designed for remote data acquisition and control, particularly in environments where the control system needs to be dispersed over a wide geographical area, while PLC is designed for real-time control and automation of industrial processes and machinery within a localized environment. RTU Communication typically supports various long-distance communication methods, including cellular, satellite, radio, and Ethernet. On contrast, PLC Communication primarily uses Ethernet and serial communications (RS-232, RS-485) within a localized network.
  3. 3.
    Programmable Logic Controllers (PLC): PLCs are specialized industrial computers used to automate and control machinery and processes. They are designed to withstand harsh industrial environments and provide reliable operation over extended periods. Key components of PLCs are:
    1. (a)
      Central Processing Unit (CPU): The CPU is the brain of the PLC, responsible for executing control instructions stored in the memory. It processes input signals, executes the control program, and sends output signals to actuators.
    2. (b)
      Input/Output (I/O) Modules: I/O modules interface the PLC with external devices. Input modules receive signals from sensors and switches, while output modules send control signals to actuators, such as motors and valves.
    3. (c)
      Communication ports : Communication ports are used for connecting to other devices and networks. Commonly used communication protocols include Ethernet, Modbus, Profibus, and DeviceNet.
    4. (d)
      Memory : Various type of memory is used in this device, it includes RAM (Random Access Memory)- a Volatile memory used for temporary data storage, ROM (Read-Only Memory)- a non-volatile memory used to store the PLC’s firmware and EEPROM (Electrically Erasable Programmable Read-Only Memory)- a Non-volatile memory used to store the control program.
    PLCs are programmed using specialized languages defined by the IEC 61131-3 standard. The most common programming languages are Ladder Logic (LAD) i.e. a graphical programming language that resembles electrical relay logic diagrams. Function Block Diagram (FBD) is another graphical language that uses blocks to represent functions and data flow. Structured Text (ST) is a high-level textual programming language similar to Pascal. It is used for complex mathematical and logical operations. Instruction List (IL) is a low-level textual language similar to assembly language. It provides fine-grained control over the PLC’s operations. Sequential Function Chart (SFC) is also a graphical language used for programming sequential processes. It represents the control process as a series of steps and transitions. The PLC operates in a cyclic manner, continuously executing the following steps Input Scan (reading input ports), Program Execution ( Executing the control logic on input data) and Output (Sending output to a output port). The real time control is achieved by executing the scan cycle rapidly, typically in milliseconds. The input interface generally receive signals from devices like push buttons, limit switches, and proximity sensors in either digital or analog signals. The output interface provide control signal to devices like relays, solenoids, and indicator lights (in digital form) or control valves, variable frequency drives (VFDs), and actuator (in analog form) etc.
    The interfaces that are generally available to communication interface with other PLC are Ethernet/IP, USB, Modbus RTU or similar protocol. A serial communication protocol used over RS-232 or RS-485. Some PLC also offers Wi-Fi or Bluetooth, that can be used for wireless programming and monitoring. PLC are generally connected with a programming terminal is used for developing, uploading, and troubleshooting the control program. Common methods of Connection are Serial Communication (RS-232/RS-485), Ethernet or USB. Some PLC also uses optical isolation using optocoupler, which consists of an LED (Light Emitting Diode) and a photodetector (phototransistor, photodiode, or photodarlington) housed in a single package. It provides a high degree of electrical isolation between different parts of the PLC system. This prevents high voltage levels or electrical noise from one part of the system from affecting other parts, and its transients. Industrial environments are typically noisy, with electromagnetic interference (EMI) and radio frequency interference (RFI) from motors, relays, and other equipment. Optical isolation helps to block these interference, ensuring that the signals within the PLC remain clean and accurate. Ground loops can occur when there are multiple ground points at different potentials, leading to unwanted currents that can interfere with signal integrity. Optical isolation breaks the electrical path, eliminating the possibility of ground loops and maintaining signal fidelity.
  4. 4.
    Human-Machine Interfaces (HMI): A Human-Machine Interface (HMI) is a sophisticated control system component that facilitates interaction between human operators and industrial control systems, such as PLCs and SCADA systems. At its core, an HMI comprises both hardware and software components designed to translate complex system data into intuitive graphical representations, allowing operators to monitor, send control signal, and optimize industrial processes effectively. Modern HMIs employ advanced visualization techniques, including real-time data graphs, schematics, and multi-layered interfaces, to enhance situational awareness and decision-making. They support various communication protocols, such as Modbus, OPC UA, and Ethernet/IP, ensuring seamless integration with diverse industrial devices and networks. Furthermore, HMIs incorporate robust security measures, including user authentication, encryption, and anomaly detection, to safeguard against cyber threats and unauthorized access. The system’s architecture often leverages embedded computing platforms, real-time operating systems (RTOS), and custom application development environments to achieve high reliability and performance in demanding industrial environments.
  5. 5.
    Data historian : A data historian is a specialized software system used in industrial environments to collect, store, and analyze time-series data from various sources. A data historian is designed to handle the massive amounts of data generated by industrial processes, often in real-time. These databases are optimized for handling time-series data, use advanced compression techniques, efficient indexing mechanisms to quickly retrieve data based on time ranges, tags, or other attributes and incorporate long-term storage strategies to archive older data, balancing between quick access for recent data and efficient storage for historical data. A typical components in the data template stored in a historian is : Tag/Point Name- a unique identifier for each data point or sensor (example: Pressure_Valve_A1), Timestamp - date and time when the data was recorded (example: 2024-07-24T14:23:00.123Z), Value- the actual measurement or value recorded by the sensor at the given timestamp, Quality/Status- indicates the quality or validity of the recorded data. This can be a binary indicator (good/bad), or a more detailed status code indicating specific conditions such as sensor error, communication failure, or manual entry (example- Good, Bad, sensor_error), Unit of Measurement- unit in which the value is measured (example- psi, $m^{3}/sec$ ). Data Source tag provide information about the origin of the data, such as the specific sensor, RTU, PLC, or other devices (example- RTU_01, PLC_Station_5), Location captures the physical or logical location of the sensor or data point (example: Plant_A, Sector_4, Lat:35.6895, Long:139.6917), Alarm/Event Information tag (if applicable) logs details about any alarms or events associated with the data point (example: High_Temperature_Alarm, Priority_1, Acknowledged).
@@ 2.2 Distributed Control Systems (DCS)
Distributed Control Systems decentralize control across various subsystems to enhance reliability and performance. Common in chemical plants, oil refineries, and large-scale production facilities. Supervisory Control and Data Acquisition (SCADA) systems and Distributed Control Systems (DCS) are both integral parts of industrial automation and control systems. While they share some similarities, they are designed for different purposes and have distinct architectures and applications. SCADA is primarily designed for high-level supervisory management and control. It is used to monitor and control processes that are distributed over large geographical areas. SCADA systems are commonly used in industries such as water and wastewater management, oil and gas pipelines, electrical power distribution, and telecommunications. DCS (Distributed Control System) is primarily designed for process control within a localized area, such as a single plant or factory. DCS systems are used for continuous and complex processes where control and monitoring are needed in a confined area. They are commonly found in industries such as chemical processing, petrochemical, oil refining, and power generation.
@@ 2.2.1 Key Differences
    Geographical scope: SCADA is designed for large-scale applications with wide geographical distribution. It is ideal for systems like water distribution networks, oil and gas pipelines, and power grids. DCS on other hand is designed for localized applications within a single plant or facility. It is ideal for controlling complex industrial processes within a confined area.
@@
    Control Approach: In SCADA, supervisory control with decision-making often done at the central control room. Local controllers like RTUs and PLCs perform basic control functions and relay data to the central system. However in DCS, distributed control with decision-making is distributed across multiple controllers. Each controller can execute complex control algorithms independently, enhancing system reliability and responsiveness.
@@
    Example Application: SCADA are mainly used in water and wastewater management, oil and gas pipelines, electrical power distribution, telecommunication systems. DCS is maily deployed in chemical processing plants, petrochemical and oil refineries, power generation plants, food and beverage processing.
@@ 2.3 Communication protocols in OT devices
    BACnet (Building Automation and Control Network): It is Open protocol designed for building automation and control systems, supports various types of data transmission (e.g., alarm, event, trend data). It is mainly used in HVAC, lighting, security, and fire detection systems in building automation.
@@ 2.4 An example Industry model with SCADA and its components
The figure 1 illustrates a mixing processing plant controlled through an Operational Technology (OT) network, highlighting the integration of Programmable Logic Controllers (PLCs), Remote Terminal Units (RTUs), and SCADA (Supervisory Control and Data Acquisition) systems. The process begins with the monitoring of critical parameters, such as temperature and RPM, via sensors attached to components like the furnace and centrifuge pump. These values are fed into the PLCs, which manage the control systems, including voltage control and motor operations, ensuring optimal function. The mixing vessel, where items are combined, is controlled by the motor and further monitored through various sensors. The entire system is supervised and controlled through SCADA stations, which communicate with the RTUs and collect historical data for future analysis. Electrical power distribution is shown in red, while OT-based communication flows are represented in green.
Figure 1: A example mixing processing plant with the OT network and PLC based controller with SCADA station.
@@ 3.1 IT and OT overlap: An attack path
Figure 2: An example IT-OT network. Here zone1 is IT network,Zone2 act as a bridge between IT and OT network, zone3 is the OT network.
By Information Technology (IT), we refer to systems used for data-centric computing. Examples are Computers, servers, enterprise applications, networking hardware. On the other hand, Operational Technology (OT) includes hardware and software that detects or causes changes through direct monitoring and control of physical devices, processes, and events. Examples are Industrial control systems (ICS), SCADA systems, PLCs, sensors etc. Modern industrial environments increasingly integrate IT systems with OT systems. Although IT/OT convergence leads to improved monitoring, data collection, and automation of industrial processes , it raises the cyber security concerns in OT networks. There are various attack vectors. Some of them are listed as below:
@@
    Malware injection: Once an IT node is compromised , a lateral movement is conducted to increase the foothold in the notwork, with aim to gain access or infiltrate to the target OT system. Then, Malware specifically designed to target OT systems is planted. Examples are Stuxnet worm, which targeted SCADA systems via infected USB drives.
@@ 3.2.1 Security Concerns in IT networks
    Access Control: It refers to implementing strict access controls to ensure that only authorized personnel can access the system. In practice, this approach emphasize on enforcement of strict access controls and least privilege principles for users with access to both IT and OT systems. It is also recommended to regularly review and audit access rights. Linux uses a traditional UNIX-like permission model that includes read (r), write (w), and execute (x) permissions for three categories of users: the file owner, the group, and others. Access Control Lists (ACLs) is another mechanism that provide a more flexible permission mechanism by allowing you to set permissions for individual users or groups beyond the owner, group, and others. Identity and Access Management (IAM) tools like Okta, Microsoft Azure Active Directory, Role-325 Based Access Control (RBAC) Implemented through Active Directory are also widely used.
@@
    Monitoring and Detection: - Deploy advanced monitoring and intrusion detection systems (IDS) to identify suspicious activities in real-time. - Use Security Information and Event Management (SIEM) systems to correlate events from IT and OT systems. Deploying Intrusion Detection Systems (IDS) and Intrusion Prevention Systems (IPS) to monitor network traffic for suspicious activity are essential components of network security. IDS is designed to monitor network traffic and alert administrators about potential security breaches, attacks, or policy violations. IPS not only detects but also prevents potential security breaches by taking action based on pre-defined rules. Both can work either on network-based or host-based. Some famous open-source IDS/IPS are Snort, Suricata, OSSEC, Bro/Zeek. Other tools that can be helpful are - Log Management Tools like Graylog (open-source log management tool that provides real-time log analysis and monitoring with features like Centralized log management, real-time search, customizable dashboards, alerting) , Forensic Analysis Tools (Autopsy-An open-source digital forensics platform for analyzing hard drives and smartphones with features like File system analysis, timeline analysis, keyword search, multimedia analysis) are all deployed and useful. SIEM (Security Information and Event Management) Tools (like Splunk with features like Event correlation, log management, compliance reporting etc), Endpoint Detection and Response (EDR) Tools (like CrowdStrike Falcon, Microsoft Defender with festures like Endpoint monitoring, behavioral analysis, automated investigation and response etc.) , Network Traffic Analysis Tools ( like wireshark, zeek with features like Deep packet inspection, real-time capture, customizable filters, extensive protocol support, event-driven scripting ).

### v2

# Cyber security of OT networks: A tutorial and survey

@@ Abstract
Cyber security, Operational Technology, OT, SCADA, PLC, OT in Healthcare.
@@ 1 Introduction
Operational Technology (OT) networks constitute critical infrastructure systems that control physical processes across sectors such as manufacturing, energy, transportation, and utilities. These networks integrate hardware and software components—including Supervisory Control and Data Acquisition (SCADA) systems, Programmable Logic Controllers (PLCs), and Distributed Control Systems (DCS)—to enable safe and efficient industrial operations. The convergence of OT and Information Technology (IT) systems, driven by digitalization and Industry 4.0 initiatives, has substantially expanded the cyber-attack surface of industrial environments [negi2024towards, knapp2024industrial, zaid2024emerging]. In contrast to conventional IT environments, OT systems exhibit unique cybersecurity challenges due to their reliance on legacy technologies, stringent real-time performance requirements, and the significant operational and financial implications of downtime. Consequently, OT cybersecurity emphasizes the protection of critical infrastructure against threats such as malware, ransomware, and state-sponsored intrusions, while preserving system availability, safety, and reliability [kalinaki2025ransomware, daniel2024emerging]. This manuscript provides a comprehensive examination of OT network architectures and their security challenges. It investigates the implications of IT–OT convergence, analyzes attack vectors and recent OT-specific cyber incidents, and reviews contemporary defensive mechanisms and tools. Furthermore, it discusses relevant regulatory frameworks, highlights emerging technological trends in OT security, and evaluates their quantitative impact on industrial resilience.
@@ 2 OT and its components
Operational Technology (OT) refers to hardware and software systems used for actuation, sensing and monitoring, computing and control of physical devices, processes, and industrial operations. In this manuscript, the context of OT is in critical sectors including manufacturing, energy, utilities, transportation, and healthcare. The key components and systems in modern OT systems (also called Industrial Control System ICS) are as below:
@@ 2.1 Supervisory Control and Data Acquisition (SCADA)
Supervisory Control and Data Acquisition (SCADA) systems provide centralized monitoring and high-level control of distributed industrial processes. They collect real-time operational data from geographically dispersed field devices and transmit it to a central control facility, where human operators can supervise, visualize, and, when required, intervene in the process. SCADA systems are extensively deployed in sectors such as power generation and distribution, oil and gas, water and wastewater management, transportation, and manufacturing [waqas2024smart].
System Overview. SCADA integrates sensing, computation, and communication layers for end-to-end process monitoring. The sensing layer captures process variables through field instruments, the control layer executes automation logic via programmable controllers, and the supervisory layer performs centralized data processing, visualization, and archiving. This architecture allows continuous process monitoring, rapid fault detection, and optimization of industrial operations across wide geographic regions.
Core Components. The main components of a SCADA system include field instruments and sensors, Remote Terminal Units (RTUs), Programmable Logic Controllers (PLCs), communication infrastructure, a master station with Human–Machine Interface (HMI), and a data historian.
Field instruments and sensors measure parameters such as temperature, pressure, flow rate, or rotational speed and generate continuous feedback for process control. Remote Terminal Units (RTUs) are microprocessor-based devices that interface with sensors, digitize analog signals, and relay data to the control center via long-distance communication channels including cellular, radio, satellite, or Ethernet links [misbahuddin2010fault]. Programmable Logic Controllers (PLCs) are industrial digital controllers responsible for executing deterministic logic in real time. They operate cyclically—scanning inputs, running control logic, and updating outputs within milliseconds—to ensure reliable process actuation. PLCs are typically programmed according to the IEC 61131-3 standard using languages such as Ladder Logic, Function Block Diagram, or Structured Text. Communication infrastructure forms the backbone interconnecting RTUs, PLCs, and the supervisory layer. Depending on the application, it may employ wired Ethernet, serial interfaces (RS-232/RS-485), or wireless networks. The SCADA master station serves as the central supervisory node that aggregates data, issues commands, and presents system status through the HMI. Finally, the data historian archives time-stamped process information for trend analysis, reporting, and compliance auditing. Typical data records include tag identifiers, timestamps, measured values, quality indicators, and event or alarm metadata.
Operational Workflow. In operation, field sensors capture process data, which RTUs and PLCs preprocess and forward through the communication network to the master station. Operators monitor key indicators on HMI displays and may dispatch control commands or re-parameterize local controllers as needed. The historian continuously logs data for performance analysis and predictive maintenance. This closed-loop interaction enables both automated and human-supervised control of industrial assets while maintaining safety and availability.
Further implementation details—such as PLC hardware architecture, memory organization, communication interfaces, and electrical isolation mechanisms—are provided in Appendix A for reference.
@@ 2.2 Distributed Control Systems (DCS)
Distributed Control Systems (DCS) decentralize control across subsystems to enhance reliability and performance and are common in chemical plants, oil refineries, and large-scale production facilities. Supervisory Control and Data Acquisition (SCADA) systems and Distributed Control Systems (DCS) are both integral parts of industrial automation and control systems. While they share some similarities, they are designed for different purposes and have distinct architectures and applications. SCADA is primarily designed for high-level supervisory management and control. It is used to monitor and control processes that are distributed over large geographical areas. SCADA systems are commonly used in industries such as water and wastewater management, oil and gas pipelines, electrical power distribution, and telecommunications. DCS (Distributed Control System) is primarily designed for process control within a localized area, such as a single plant or factory. DCS systems are used for continuous and complex processes where control and monitoring are needed in a confined area. They are commonly found in industries such as chemical processing, petrochemical, oil refining, and power generation.
@@ 2.2.1 Key Differences
    Geographical scope: SCADA is designed for large-scale applications with wide geographical distribution. It is ideal for systems like water distribution networks, oil and gas pipelines, and power grids. DCS, on other hand, is designed for localized applications within a single plant or facility. It is ideal for controlling complex industrial processes within a confined area.
@@
    Control Approach: In SCADA, supervisory control with decision-making often done at the central control room. Local controllers like RTUs and PLCs perform basic control functions and relay data to the central system. However, in DCS, distributed control with decision-making is distributed across multiple controllers. Each controller can execute complex control algorithms independently, enhancing system reliability and responsiveness.
@@
    Example Application: SCADA is mainly used in water and wastewater management, oil and gas pipelines, electrical power distribution, telecommunication systems. DCS is mainly deployed in chemical processing plants, petrochemical and oil refineries, power generation plants, food and beverage processing.
@@ 2.3 Communication protocols in OT devices
    BACnet (Building Automation and Control Network): It is an open protocol designed for building automation and control systems, supports various types of data transmission (e.g., alarm, event, trend data). It is mainly used in HVAC, lighting, security, and fire detection systems in building automation.
@@ 2.4 An example Industry model with SCADA and its components
The figure 1 illustrates a mixing process plant controlled through an Operational Technology (OT) network, highlighting the integration of Programmable Logic Controllers (PLCs), Remote Terminal Units (RTUs), and SCADA (Supervisory Control and Data Acquisition) systems. The process begins with the monitoring of critical parameters, such as temperature and RPM, via sensors attached to components like the furnace and centrifuge pump. These values are fed into the PLCs, which manage the control systems, including voltage control and motor operations, ensuring optimal function. The mixing vessel, where items are combined, is controlled by the motor and further monitored through various sensors. The entire system is supervised and controlled through SCADA stations, which communicate with the RTUs and collect historical data for future analysis. Electrical power distribution is shown in red, while OT-based communication flows are represented in green.
Figure 1: An example mixing processing plant with the OT network and PLC based controller with SCADA station.
@@ 3.1 IT and OT overlap: An attack path
Figure 2: Example IT–OT network architecture. Zone 1 represents the IT network, Zone 2 is the IT–OT bridge, and Zone 3 represents the OT network.
By Information Technology (IT), we refer to systems used for data-centric computing. Examples are Computers, servers, enterprise applications, networking hardware. On the other hand, Operational Technology (OT) includes hardware and software that detects or causes changes through direct monitoring and control of physical devices, processes, and events. Examples are Industrial control systems (ICS), SCADA systems, PLCs, sensors etc. Modern industrial environments increasingly integrate IT systems with OT systems. Although IT/OT convergence leads to improved monitoring, data collection, and automation of industrial processes it raises the cyber security concerns in OT networks. There are various attack vectors. Some of them are listed as below:
@@
    Malware injection: Once an IT node is compromised, attackers move laterally to gain a foothold in the OT network. Then, malware specifically designed to target OT systems is planted. Examples are Stuxnet worm, which targeted SCADA systems via infected USB drives.
@@ 3.2.1 Security Concerns in IT networks
    Access Control: It refers to implementing strict access controls to ensure that only authorized personnel can access the system. In practice, this approach emphasizes enforcement of strict access controls and least privilege principles for users with access to both IT and OT systems. It is also recommended to regularly review and audit access rights. Linux uses a traditional UNIX-like permission model that includes read (r), write (w), and execute (x) permissions for three categories of users: the file owner, the group, and others. Access Control Lists (ACLs) is another mechanism that provide a more flexible permission mechanism by allowing you to set permissions for individual users or groups beyond the owner, group, and others. Identity and Access Management (IAM) tools like Okta, Microsoft Azure Active Directory, Role-325 Based Access Control (RBAC) Implemented through Active Directory are also widely used.
@@
    Monitoring and Detection: - Deploy advanced monitoring and intrusion detection systems (IDS) to identify suspicious activities in real-time. - Use Security Information and Event Management (SIEM) systems to correlate events from IT and OT systems. Deploying Intrusion Detection Systems (IDS) and Intrusion Prevention Systems (IPS) to monitor network traffic for suspicious activity are essential components of network security. IDS is designed to monitor network traffic and alert administrators about potential security breaches, attacks, or policy violations. IPS not only detects but also prevents potential security breaches by taking action based on pre-defined rules. Both can work either on network-based or host-based. Some famous open-source IDS/IPS are Snort, Suricata, OSSEC, Bro/Zeek. Other tools that can be helpful are - Log Management Tools like Graylog (open-source log management tool that provides real-time log analysis and monitoring with features like Centralized log management, real-time search, customizable dashboards, alerting) , Forensic Analysis Tools (Autopsy-An open-source digital forensics platform for analyzing hard drives and smartphones with features like File system analysis, timeline analysis, keyword search, multimedia analysis) are all deployed and useful. SIEM (Security Information and Event Management) Tools (like Splunk with features like Event correlation, log management, compliance reporting etc), Endpoint Detection and Response (EDR) Tools (like CrowdStrike Falcon, Microsoft Defender with features like Endpoint monitoring, behavioral analysis, automated investigation and response etc.) , Network Traffic Analysis Tools ( like wireshark, zeek with features like Deep packet inspection, real-time capture, customizable filters, extensive protocol support, event-driven scripting ).